Form an external point P tangents are drawn to the parabola y^2=4ax find the equation to the locus P when these tangents make an angle theta1 and theta2 with the axis, such that (costheta1)(costheta2) is constant.

Grade:12

1 Answers

Dilli PRASAD
5 years ago
Cost12cost22=k2
Sect12 sect22=1/k2
(1+tant12)(1+tant22)=1/k2
(1+m12)(1+m22)=1/k2
Afteter multiplication put m1+m2=y/x
m1.m2=a/x